OpenStudy (anonymous):

find the indicated outputs for f(x)=5xwith small 2 above - 2x f(0)= f(-1) = f(2)=

OpenStudy (anonymous):

????? cud not understand the given function....

OpenStudy (anonymous):

i solved this way: f(0) = 0 f(-1) = 5xwith small 2 -2x f(-1)=5(-1)small2 -2(-1) =5+2 =7 f(2)=5(2)small 2 -2(2) =20-4 =16

OpenStudy (anonymous):

so my answers are 0 7 16

OpenStudy (anonymous):

pls use the equation box to type the question properly

OpenStudy (anonymous):

where is equation box i'm new at this

OpenStudy (anonymous):

is it \[5x ^{2}/-2x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look below the box in which you type answer. just below it u hv "Equation" and "Attach File" buttons side- be-side on the left...

OpenStudy (anonymous):

k thnks

OpenStudy (anonymous):

wait i am checking yr answers.......

OpenStudy (anonymous):

except for first answers, other r wrong.. I'll show how it is to be done...

OpenStudy (anonymous):

^2 means "raised to power 2" 5x^2 f(x) = -------- -2x 5 (0)^2 5 * 0 0 f(0) = -------- = ------ = ----- = 0 -2(0) 0 0

OpenStudy (anonymous):

5 (-1)^2 5 * 1 5 f(-1) = -------- = ------ = ----- -2(-1) 2 2 and 5 (2)^2 5 * 4 20 f(2) = -------- = ------ = ----- = -5 -2(2) -4 -4

OpenStudy (anonymous):

is it clear now how to do such sums???

OpenStudy (anonymous):

yes it is now I had those other two answers but for some reason I did not stop there thank you and have a good morning

OpenStudy (anonymous):

good morning to u too!!

OpenStudy (anonymous):

however u were doing them wrong as for second u were getting 5+2 instead of 5/2

OpenStudy (anonymous):

for third u were getting 20-4 =16 instead of 20/-4 = -5

OpenStudy (anonymous):

when i worked them earlier I had 0, 5/2 and -5

OpenStudy (anonymous):

anyway thanks for putting me back on track

OpenStudy (anonymous):

given \[5x^2/-2x\] you should simplify first to -5x/2 and then input your numbers

OpenStudy (anonymous):

yes that is an excellent suggestion by nobody..... medal for him.....